OpenStudy (anonymous):

x-2y=-4 y=1/2x+2 What is the solution and number of solutions?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Is this two problems?

OpenStudy (anonymous):

x-2(1/2x+2)=-4 x-5x=-4 [-4x=-4]1/4 x=1 maybe, hehehe

OpenStudy (anonymous):

and what would be the number of solutions, and the solution?

OpenStudy (anonymous):

@bahupepe how did you get -5x on the second line? o.O